the height to which a cylindrical vessel be filled with a homogenous liquid ,to make the average force with which the liquid presses the side of the vessel equal to the force exerted by the liquid on the bottom of the vessel ,equal to -?

Grade:11

3 Answers

Arun
6 years ago
Dear Abhinav
 
Pressure = \rho* h* g
and pressure on vertical surface = (1/2) \rho* h* g
Now according to the problem
 
Force at the bottom = force on the vertical surface
\rhohg * \pi* r²  = (1/2)\rhohg * 2*\pi* r*h
h = r
 
Hence height will be equal to the radius of vessel.
 
Regards
Arun (askIITians forum expert)
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Pressure = ρ* h* g
and pressure on vertical surface = (1/2) ρ* h* g
 
Now according to the problem
Force at the bottom = force on the vertical surface
ρhg * π* r²  = (1/2) ρ hg * 2* π* r*h
Hence, h = r
 
Hence height will be equal to the radius of vessel.
 
Thanks and regards,
Kushagra
pushkar
3 years ago
Bhai mera question bhi koi solve kro
 
..................,......................................
 
...........................
.......